OpenStudy (anonymous):

There is a triangle ABC, comparison between the tan A: tan B: tan C = 1:2:3. How to find AB:AC ?

OpenStudy (blockcolder):

In a triangle, tan A + tan B + tan C = tan A tan B tan C.

OpenStudy (p0sitr0n):

sinA/cosA : sinB/cosB : sinC:cosC

OpenStudy (anonymous):

@blockcolder : then??

OpenStudy (blockcolder):

Let tan A=k, tan B=2k, tan C=3k. From my previous reply, we have k+2k+3k=k*2k*3k 6k=6k^3 6k^3-6k=0 6k(k^2-1)=0 k=-1, 0, 1 k=0 would mean tan A=0 and A=0, so this is impossible. k=-1 would mean tan A=-1 and A=-pi/4 so again this is impossible. k=1 would mean tan A=1 and A=pi/4.

OpenStudy (blockcolder):



OpenStudy (anonymous):

then tan B=2. so B is 63,43. then tan C=3 so C is 71.565 is it correct?

OpenStudy (blockcolder):

Now use the law of sines to find the ration AB:AC.